Determine the exact number of square inches of pizza I will get from 2- 7” pizzas and then from a 14”

pizza, letting me know the outcome of your computations.

Then, determine how many more square inches of pizza I will get if I order 1- 14” pizza, instead of 1- 7” pizza, again,

letting me know the outcome of your computations.

Also, if I order the 14” pizza, beside myself, I will have to share it equally with 5 friends.

Exactly how much pizza will I get to eat?
Answer #1
  • Square inches of pizza get from two 7" Pizzaa

                  Pizza is circle,so we have to find square inches by finding Pi*r2 where Pi=3.14 and r=given inch/2

                    Here, 3.14*(7/2)2 = 38 square inches for 1 pizza

                      so, for 2 pizzas= 76 square inches

  • Square inches of pizza get from two 14" Pizzaa

                  Pizza is circle,so we have to find square inches by finding Pi*r2 where Pi=3.14 and r=given inch/2

                    Here, 3.14*(14/2)2 = 154 square inches for 1 pizza

                      so, for 2 pizzas= 308 square inches

  • Determine how many more square inches of pizza I will get if I order 1- 14” pizza, instead of 1- 7” pizza

                       1-14" pizza=154 square inches

                       1-7 pizza=38 square inches

                       Difference=154-38=116 square inches

  • I order the 14” pizza, beside myself, I will have to share it equally with 5 friends.

    Exactly how much pizza will I get to eat?

                14'' pizza = 154 square inches

                Which divided into 6, then 154/6=26 inches

                  I get to eat is 26 inches